# Theory of intervalley-coherent AFM order and topological superconductivity in tWSe2

### v1

@@ VI Evolution of the SC phase in the intermediate twist angle regime
Finally, we map out phase diagrams in the intermediate twist angle regime $\theta=\{4.1^{\circ},4.5^{\circ}\}$ as summarized in Fig. 6 (a,b). Decreasing the twist angle quenches the kinetic energy scales such that the moiré bands react more sensitively to the external displacement field. As a result, the position of the HOVHS is moved to smaller values of $E_{z}$, see blue circle in Fig. 5 (c,d) and SM SM . As the ratio between the Wannier function spread and the gate screening distance $\Omega/\xi$ decreases with the twist angle, long-ranged density-density interaction between the Wannier orbitals are successively suppressed and we therefore resort to the Hubbard Hamiltonian defined in Eq. (1). The values of the Hubbard interactions for different twist angles are listed in the SM SM .
@@
Apart from the enhanced impact of the displacement field on the single-particle bandstructure, decreasing the twist angle increases the ratio $U/W$ and places the system in a moderate interacting regime. As a consequence, the emergent IVC-AFM state and the chiral $d/p$-wave SC state manifest at larger scales $\Lambda_{c}$ indicating a higher critical temperature of respective phases. Consistent with the analysis brought forward in Section IV.2, this enhances the asymmetry of the IVC-AFM with respect to the van-Hove line as demonstrated in Fig. 6 (a,b). At $\theta=4.5^{\circ}$, the maximal critical scale $\Lambda_{c}$ of the IVC-AFM phase is found around $E_{z}\sim 15\,\mathrm{meV}$ and is clearly displaced from the van-Hove line, where instead regions of chiral $d/p$-wave SC prevail. At $\theta=4.1^{\circ}$ the IVC-AFM state is displaced to the vicinity of $n/n_{0}\sim 1$ and the superconducting region is shrunk in size; a trend that is observed in small twist angle tWSe2 in Ref. Xia et al. (2024). In particular, superconductivity does not extend to $E_{z}=0\,\mathrm{meV}$, but is tied to an intermediate displacement field regime $5\,\mathrm{meV}<E_{z}<10\,\mathrm{meV}$.
VII Conclusion and outlook
@@
In this work, we provide a complete and unbiased characterization of correlated states in $\theta=4^{\circ}\dots 5^{\circ}$ tWSe2 from first principles by studying the impact of gate-screened Coulomb interactions with state-of-the-art functional renormalization group techniques. Our analysis reveals that inter-valley coherent anti-ferromagnetic order (IVC-AFM) and chiral, mixed-parity $d/p$-wave superconductivity collaborate along the gate-tunable van-Hove line. The superconducting pairing glue for the unconventional SC state is provided by spin-fluctuations arising in regions of incommensurate IVC-AFM order below the HOVHS, where the Fermi surface pockets around the valleys $K^{\nu}$ are broadened and therefore weaken scattering between opposite spin-valley sectors. We attribute the asymmetry of particle-hole instabilities observed in experiment Xia et al. (2024); Guo et al. (2024) to the formation of hole pockets around $K^{\nu}$ that are broadened at larger scales (temperature) such that the dominant ordering vector locks back to $K^{\nu}$ at positions that are offset from the van-Hove line  [^1] We foresee this scale-dependent nesting as a general feature of weak-to-moderately interacting fermions with HOVHS.. As the hole pockets are equally polarized on the $T/H_{2}$-orbital, this stresses the importance of describing the low-energy physics of tWSe2 in the multi-orbital setup presented in this manuscript. One aspect of the experiment of Guo et. al. Guo et al. (2024) that is however not clearly reproduced by our calculations is the non-monotonic variation of antiferromagnetism with carrier concentration at fixed displacement field. At displacement fields slightly greater than those where superconductivity is observed, experiment finds evidence for antiferromagnetism at carrier concentrations slightly lower than and slightly higher than the van-Hove concentrations, but not when the carrier density is tuned to the van-Hove concentration. Understanding this is behavior is an important topic for future research. Further, we demonstrate that long-ranged density-density interactions do not alter the phase diagram of tWSe2 qualitatively, but change the phase boundaries and critical temperature due to the coupling of the IVC-AFM order parameter to a spin/valley bond order (SBO) that acts as an interaction-induced spin-orbit coupling term. We therefore argue that the presence of longer-ranged density-density interactions will enhance the critical temperature $T_{c}$ of the SC phase, however, decrease its overall size in the phase diagram due to the competition with the IVC-AFM state.
@@
Finally, we show that as the twist angle is decreased to $\theta=4.1^{\circ}$ the regions of SC order are displaced to smaller values of the external displacement field ($5\,\mathrm{meV}<E_{z}<10\,\mathrm{meV}$) and to hole densities closer to $n/n_{0}\sim 1$ in good agreement with experimental measurements. It is interesting to consider these results in light of the finding of Xia et. al. Xia et al. (2024) that in a $\theta=3.65^{\circ}$ angle device, a superconducting region exists very near $n/n_{0}=1$ and vanishing displacement field, giving way to an insulating phase as the displacement field is increased. Whether this superconductor-insulator transition is a consequence of strong coupling Mott/spin liquid physics or can be understood within the weak-to-moderate coupling physics captured by the FRG approach used here is an important open question.
The validity of the static four-point FRG employed in this work is bound to the weak to moderate interacting regime $U/W<1$ that is relevant for the twist angles considered within this work. By further decreasing the twist angle, we expect that the system is eventually tuned into a strongly interacting regime. It is thus a highly anticipated avenue of future research to treat this regime with state-of-the art methods such as, e.g. DMFT or slave Bosons to learn about the properties of strongly correlated and ordered states in tWSe2.
@@ Acknowledgements.
This work was supported by the Excellence Initiative of the German federal and state governments, the Ministry of Innovation of North Rhine-Westphalia and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). AF and DMK acknowledge funding by the DFG within the Priority Program SPP 2244 “2DMP” – 443274199. LK acknowledges support from the DFG through Project-ID 258499086 – SFB 1170 and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat, Project-ID 390858490 – EXC 2147. LK and TOW greatfully acknowledge support from the DFG through FOR 5249 (QUAST, Project No. 449872909) and SPP 2244 (Project No. 422707584). SR is supported by the DFG research unit FOR 5242 (WE 5342/7-1, project No. 449119662). TOW is supported by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the DFG (EXC 2056, Project ID 390715994). AR acknowledges support by the European Research Council (ERC-2015-AdG694097), the Cluster of Excellence ‘Advanced Imaging of Matter’ (AIM), Grupos Consolidados (IT1249-19) and Deutsche Forschungsgemeinschaft (DFG) – SFB-925 – project 170620586. LX acknowledges supported by the National Key Research and Development Program of China (Grant No. 2022YFA1403501), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2022B1515120020), the National Natural Science Foundation of China (Grant No. 62341404), Hangzhou Tsientang Education Foundation and the Max Planck Partner group programme. AJM acknowledges support from Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award (DE-SC0019443). DMK and AR acknowledge support by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. Computations were performed on the HPC system Ada at the Max Planck Computing and Data Facility. The Flatiron Insitute is a division of the Simons Foundation.

### v2

@@ VIII Evolution of the IVC-AFM order and superconductivity across twist angles
After having established our methodology for $\theta=5.08^{\circ}$ tWSe2, we map out the FRG phase diagrams in the twist angle regime $\theta=\{3.7^{\circ},3.9^{\circ},4.1^{\circ},4.5^{\circ}\}$ as summarized in Fig.˜9 (a-d). Decreasing the twist angle quenches the kinetic energy scales such that the moiré bands react more sensitively to the external displacement field. As a result, the position of the HOVHS is moved to smaller values of $E_{z}$, see blue circle in Fig.˜9 (a-d). As the ratio between the Wannier function spread and the gate screening distance $\Omega/\xi$ decreases with the twist angle, long-ranged density-density interaction between the Wannier orbitals are successively suppressed and we therefore resort to the Hubbard Hamiltonian defined in Eq. (2). We provide a detailed overview of the twist angle dependence of the DOS and the values of the density-density interaction tensor in the SM [76]. At small twist angles, we only perform FRG simulations for experimentally relevant values of the displacement field corresponding (black dashed line).
@@
Decreasing the twist angle further increases the ratio $U/W$ and places the system in a moderate-to-strongly interacting regime. As a consequence, the emergent IVC-AFM order manifests at larger scales $\Lambda_{c}$ indicating a higher critical temperature of magnetic ordering. Consistent with the analysis brought forward in Section˜V.1, this enhances the asymmetry of the IVC-AFM order with respect to the van-Hove line as demonstrated in Fig.˜9 (a-d). At $\theta=4.5^{\circ}$, the maximal critical scale $\Lambda_{c}$ of the IVC-AFM phase is found around $E_{z}\sim 18\,\mathrm{meV}$ and is clearly displaced from the van-Hove line. At $\theta=4.1^{\circ}$, the IVC-AFM order extend towards the commensurate hole filling of $n/n_{0}\sim-1$ and replaces former superconducting regions along the van-Hove line. At half-filling $n/n_{0}\sim-1$, quasiparticle renormalizations in the IVC-AFM phase can fully gap the Fermi surface, opposite to the scenario at larger twist angles where IVC-AFM order at incommensurate densities $n/n_{0}\neq-1$ reconstructs the Fermi surface and the system is a partially-gapped antiferromagnetic metal, c.f. Fig. 4 (a). The transition from a partially-gapped IVC-AFM metal to a fully-gapped insulating state in the intermediate twist angle regime $\theta\sim 4.5^{\circ}$ as predicted by our theory is in line with previous experimental measurements [37, 39, 38].
The asymmetric bending of IVC-AFM domains is further enhanced for the smaller twist angles $\theta=3.9^{\circ}$ and $\theta=3.7^{\circ}$ such that SC order is eventually completely displaced from the van-Hove line as shown in Fig. 9 (c,d). At small twist angles and small values of the displacement field $E_{z}<8$ meV, additional magnetic phases emerge in the vicinity of the van-Hove line. To analyze these states, we provide a full characterization of the leading FRG eigenvector $\phi^{L}_{\kappa}(\bm{q}_{C})$ in the crossed particle-hole channel ($C$) by color-coding the evolution of the leading momentum transfer $\bm{q}_{C}$ in the irreducible wedge of the moiré Brillouin zone. We further distinguish between IVC order $\phi^{L}_{\kappa}(\bm{q}_{C})\sim\sigma_{x,y}$] or valley-polarized (VP) $\phi^{L}_{\kappa}(\bm{q}_{C})\sim\sigma_{z}$] solutions as shown in the middle row of Fig. 9. Along the van-Hove line and for small values of the displacement field, $\theta=3.7^{\circ}$ tWSe2 features domains of ferromagnetic VP order ($\bm{q}_{C}=\Gamma$, blue), albeit their critical temperature is lower compared to the IVC-AFM domains. In their vicinity, smaller domains of striped-IVC order ($\bm{q}_{C}=M$, red) emerge. For larger values of the displacement field, these domains smoothly transition into regions of incommensurate IVC-AFM order ($\bm{q}_{C}=K\to\Gamma$, darkorange) until eventually recovering commensurate IVC-AFM order ($\bm{q}_{C}=K$, orange) that follows the van-Hove line. For larger twist angles, the additional magnetic phases are weakened and only the IVC-AFM order survives in the vicinity of the van-Hove line.
Figure 9: Evolution of IVC-AFM domains and SC order in the low-energy phase diagram of tWSe2 in the twist angle regime $\theta=3.7^{\circ}\dots 4.5^{\circ}$. (a) FRG phase diagram of $\theta=4.5^{\circ}$ tWSe2. The upper row shows the critical scale $\Lambda_{\mathrm{c}}\sim T_{\mathrm{c}}$ of the leading spin/charge density wave (DW, red) and superconducting instability (SC, blue). The DOS of the normal-state (grey) is given as reference and the blue circle indicates the position of the HOVHS. The grey-hatched area represents the experimentally inaccessible regime. The middle panel shows the leading FRG eigenvector $\phi^{L}_{\kappa}(\bm{q}_{C})$ for all DW instabilities; color-coding the evolution of the leading momentum transfer $\bm{q}_{C}$ in the irreducible wedge of the moiré Brillouin zone and differentiating between IVC $\phi^{L}_{\kappa}(\bm{q}_{C})\sim\sigma_{x,y}$] or valley-polarized (VP) $\phi^{L}_{\kappa}(\bm{q}_{C})\sim\sigma_{z}$] order. The lower panel shows the amplitude (opacity) and phase (color) of the SC order parameter on the Fermi surface (grey line) segregated into singlet (triplet) components $\psi_{0}(\bm{k})$ ($d_{z}(\bm{k})$) on the $T/H_{2}$-orbitals. (b,c,d) FRG phase diagrams for $\theta=4.1^{\circ}$, $\theta=3.9^{\circ}$ and $\theta=3.7^{\circ}$. Regions of IVC-AFM order (convex hull) evolve to smaller values of the displacement field $E_{z}$ and bend towards $n/n_{0}\sim-1$, causing a pronounced asymmetry with respect to the van-Hove line (grey). Regions of chiral $d/p$-wave SC order are shrunk in size and are eventually displaced from the van-Hove line in favor of IVC-AFM order at $\theta\sim 4.0^{\circ}$. In this regime, SC domains move towards $n/n_{0}\sim-1$ and their critical temperature is decreased due to the change in fermiology and the suppression of DOS away from the van-Hove line. Chiral $d/p$-wave SC order remains the leading superconducting instability, albeit the SC amplitude is constrained to the $T$-orbital signaling a transition from multi-orbital to single-orbital SC order.
VIII.1 Evolution of SC order towards smaller twist angles: Multi-orbital to single-orbital SC order
At $\theta\sim 3.9^{\circ}$, the superconducting region in the phase diagram is displaced from the van-Hove line and moves towards $n/n_{0}\sim-1$ and increasingly smaller values of $E_{z}$. Away from the van-Hove line, the fermiology of the system changes from the triangular Fermi pockets centered around $K^{\nu}$ to a single connected hexagon interleaved with the moiré BZ, see Fig. 9. The detuning from the van-Hove line causes a decrease in the DOS that leads to a sudden decrease of the critical scale $\Lambda_{\mathrm{c}}$ associated to the SC order parameter, albeit the system is tuned into a more strongly-correlated regime. We observe that the symmetry classification and topological properties of the superconducting state are not altered and the leading pairing instability across all twist angles remains the chiral, mixed-parity $d/p$-wave. However, while at $\theta=5.08^{\circ}$ the amplitude on the $T/H_{2}$ bonds is nearly equal, the SC order parameter at lower twist angles is only facilitated by the $T$ orbital that resides in the $MM$ stacking regions of the superlattice. Therefore, we claim that the superconducting order parameter changes from a multi-orbital to single-orbital character as the twist angle is decreased. The SC state at lower twist angles may be compatible with early theoretical works based on the single-band ‘moiré-Hubbard model’ [70], where superconductivity has been reported away from the van-Hove line at hole densities $n/n_{0}\sim-1$ [42].
IX Conclusion and outlook
@@
In this work, we provide a complete and unbiased characterization of correlated states in $\theta=3.7^{\circ}\dots 5^{\circ}$ tWSe2 from first principles by studying the impact of gate-screened Coulomb interactions with state-of-the-art functional renormalization group techniques. Our analysis reveals that inter-valley coherent anti-ferromagnetic order (IVC-AFM) and orbital-selective, chiral $d/p$-wave superconductivity collaborate along the gate-tunable van-Hove line. The superconducting pairing glue for the unconventional SC state is provided by spin-fluctuations arising in regions of incommensurate IVC-AFM order below the HOVHS, where the Fermi surface pockets around the valleys $K^{\nu}$ are broadened and therefore weaken scattering between opposite spin-valley sectors. We attribute the asymmetry of particle-hole instabilities observed in experiment [37, 39] to the formation of hole pockets around $K^{\nu}$ that are broadened at larger scales (temperature) such that the dominant ordering vector locks back to $K^{\nu}$ at positions that are offset from the van-Hove line  [^1] We foresee this scale-dependent nesting as a general feature of weak-to-moderately interacting fermions with HOVHS.. As the hole pockets are equally polarized on the $T/H_{2}$-orbital, this stresses the importance of describing the low-energy physics of tWSe2 in the multi-orbital setup presented in this manuscript. By performing a comprehensive FRG+MF analysis within the ordered IVC-AFM phase, we show that our calculations can explain pivotal aspects of the experimental phase diagram including an asymmetric occurrence of IVC-AFM domains with respect to the tunable van-Hove line, the transition from a single-peak to a double-peak structure in the density of states (DOS) and the subtle interplay between IVC-AFM order and superconductivity along the van-Hove line. Our numerical framework unveils that particle-particle screening leads to a significant suppression of the bare down-folded Coulomb interaction, which explains the overestimation of electronic order in Hartree-Fock based approaches that do not resolve vertex corrections. We further provided an in-depth analysis of spectral properties in the SC phase and unravel a strong orbital-selectivity that manifests in nodal signatures in the local density of states (LDOS) and is related to the topological properties of the Wannier model. Further, we demonstrate that long-ranged density-density interactions do not alter the phase diagram of tWSe2 qualitatively, but change the phase boundaries and critical temperature due to the coupling of the IVC-AFM order parameter to a spin/valley bond order (SBO) that acts as an interaction-induced spin-orbit coupling term. We therefore argue that the presence of longer-ranged density-density interactions will enhance the critical temperature $T_{c}$ of the SC phase, however, decrease its overall size in the phase diagram due to the competition with the IVC-AFM state.
Our theoretical predictions provide guidance for future experiments aiming to verify the nature of the correlated states in tWSe2. Besides magnetic circular dichroism (MCD) measurements [38] of the insulating state at $n/n_{0}=1$ in 4.6 tWSe2 suggestive of inter-valley coherent order, future experiments could probe the inhomogeneous spin spectral density in the IVC-AFM phase associated to the multi-orbital physics via local spectroscopy techniques like (magnetic) scanning tunneling microscope [113]. Another promising route is the newly developed quantum twisting microscope [114, 115] that could be exploited to directly probe the momentum-resolved spectral function $A(\bm{k},\omega)$ in the correlated metallic (insulating) IVC-AFM phase as provided in Fig. 4. As the chiral $d/p$-wave superconductor descends from the IVC-AFM phase and is of Ising-type, i.e. Cooper pairs are formed by electrons from different valleys with opposite spins due to strong intrinsic Ising-SOC, we expect a weak (strong) dependence of the SC state subject to in-plane (out-of-plane) magnetic fields [116, 117, 118] with possibly large Pauli-limit violations for in-plane magnetic fields.
@@
Finally, we show that as the twist angle is decreased to $\theta\sim 4.0^{\circ}$ the regions of SC order are displaced to smaller values of the external displacement field ($5\,\mathrm{meV}<E_{z}<10\,\mathrm{meV}$) and to hole densities closer to $n/n_{0}\sim-1$ in good agreement with recent experimental measurements [38]. For smaller twist angles $\theta\lesssim 4.0^{\circ}$, SC order is eventually displaced from the van-Hove line, which leads to a suppression of the critical temperature due to the reduction in DOS. It is interesting to consider these results in light of the finding of Xia et. al. [37] that in a $\theta=3.65^{\circ}$ angle device, a superconducting region exists very near $n/n_{0}=1$ and vanishing displacement field, giving way to an insulating phase as the displacement field is increased.
We further note that the quantitative validity of the static four-point FRG employed in this work is bound to the weak-to-moderate interacting regime $U/W<1$ that seems relevant for the large-to-intermediate twist angle regime considered within this work. By further decreasing the twist angle, we expect that the system is eventually tuned into a strongly interacting regime, in which dynamical correlations gain in importance [9]. One may ask, for example, whether this superconductor-insulator transition observed in the $3.65^{\circ}$ twist angle device can be understood in terms of the weak/intermediate coupling theory presented here or has properties inconsistent with our theory, requiring consideration of strong coupling Mott/spin liquid physics. It is thus a highly anticipated avenue of future research to treat this regime with state-of-the art methods such as, e.g. DMFT, DMFT+FRG [119] or slave Bosons to learn about the interplay of Mott insulators and the ordered IVC-AFM (chiral $d/p$-wave superconductor) proposed by FRG. The methodological framework developed here can further be applied to tackle related questions concerning the interplay of magnetic order and superconductivity in rhombohedral multilayer graphene [120, 121] via effective multi-orbital Wannier models as e.g. outlined in Ref. [79] or to investigate the role of iso-spin fluctuations in twisted bilayer graphene by treating electronic correlations in multi-orbital Wannier models [69, 68] with DMFT [122, 123] as starting point.
@@ Acknowledgements.
This work was supported by the Excellence Initiative of the German federal and state governments, the Ministry of Innovation of North Rhine-Westphalia and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). AF and DMK acknowledge funding by the DFG within the Priority Program SPP 2244 “2DMP” – 443274199. LK acknowledges support from the DFG through Project-ID 258499086 – SFB 1170 and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat, Project-ID 390858490 – EXC 2147. LK and TOW greatfully acknowledge support from the DFG through FOR 5249 (QUAST, Project No. 449872909) and SPP 2244 (Project No. 422707584). SR is supported by the DFG research unit FOR 5242 (WE 5342/7-1, project No. 449119662). TOW is supported by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the DFG (EXC 2056, Project ID 390715994). AR acknowledges support by the European Research Council (ERC-2015-AdG694097), the Cluster of Excellence ‘Advanced Imaging of Matter’ (AIM), Grupos Consolidados (IT1249-19) and Deutsche Forschungsgemeinschaft (DFG) – SFB-925 – project 170620586. LX acknowledges supported by the National Key Research and Development Program of China (Grant No. 2022YFA1403501), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2022B1515120020), the National Natural Science Foundation of China (Grant No. 62341404), Hangzhou Tsientang Education Foundation and the Max Planck Partner group programme. AJM acknowledges support from Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award (DE-SC0019443). DMK and AR acknowledge support by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. Computations were performed on the HPC system Ada at the Max Planck Computing and Data Facility. The Flatiron Institute is a division of the Simons Foundation.
Data availability
The data that support the findings of this article are openly available [124].